OpenStudy (anonymous):

What is the simplest polynomial function that can be written with zeros 4+1 and ± √2 ?

OpenStudy (anonymous):

some weird cut and past here

OpenStudy (anonymous):

it is maybe \(4+i\) and \(\pm\sqrt{2}\) ?

OpenStudy (anonymous):

it doesn't like you copying the math written as an object

OpenStudy (anonymous):

@satellite73 yes fancy italic i. I'm not asking anyone to do my homework. i just don't know how to figure this out at all

OpenStudy (anonymous):

so it is definitely \(4+i\) right?

OpenStudy (anonymous):

@satellite73 and ±2√

OpenStudy (anonymous):

you are going to have a big of work to do, but most of it is not too hard if the zeros are \(\pm\sqrt2\) then you know it factors as \[(x-\sqrt{2})(x+\sqrt{2})=x^2-2\]

OpenStudy (anonymous):

hmm am confused about the \(\pm2\sqrt{}\) that does make sense to me

OpenStudy (anonymous):

could it have been \(\pm\sqrt2\) instead? that i understand

OpenStudy (anonymous):

@satellite73 wow i can't type at all today. the question states and ±√2.

OpenStudy (anonymous):

oh good now we can get somewhere

OpenStudy (anonymous):

from those two zeros , you know it has the factors \[(x-\sqrt2)(x+\sqrt2)\] and when you multiply them out, you get \[x^2-2\] so far so good?

OpenStudy (anonymous):

we are not done, i am just asking if that step is okay so far

OpenStudy (anonymous):

okay, that makes sense.

OpenStudy (anonymous):

now we have to come up with a polynomial that has a zero of \(4+i\) which will also have a zero of \(4-i\) the conjugate there are many ways to do this one of the simpler ones is to set \[x=4+i\] and work backwards

OpenStudy (anonymous):

here goes \[x=4+i\] subtract \(4\) get \[x-4=i\] then square both sides to get \[(x-4)^2=i^2\] or \[x^2-8x+16=-1\] finally add \(1\) to get \[x^2-8x+17\] as your polynomial with zero \(4+i\)

OpenStudy (anonymous):

you last job is to multiply \[(x^2-2)(x^2-8x+17)\] to get your answer in standard form

OpenStudy (anonymous):

@satellite73 x^4−8x^3+15x^2+16x−34 ?

OpenStudy (anonymous):

if you have to do a lot of these it is easiest just to memorize that if \(a+bi\) is the zero of a quadratic polynomial then the polynomial is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

looks good to me wanna check it?

OpenStudy (anonymous):

@satellite73 thank you so much!

OpenStudy (anonymous):

here you can check that it has the right zeros http://www.wolframalpha.com/input/?i=x^4%E2%88%928x^3%2B15x^2%2B16x%E2%88%9234+

OpenStudy (anonymous):

yw, now i have a question

OpenStudy (anonymous):

who is "gisa" that should drop dead?

OpenStudy (anonymous):

or is that a play on "drop dead gorgeous" ?

OpenStudy (anonymous):

@satellite73 i made this account when I was going through my "Scene emo" stage and it sounded realllllly cool.. at the time.

OpenStudy (anonymous):

emo nice favorite was "if arsenic fails, try algebra"